# Learning Geodesics of Geometric Shape Deformations From Images

### v1

@@ 3.2 Joint Learning of GDN
The loss function of GDN integrates losses from both an unsupervised registration network and the geodesic neural operators informed by real numerical solutions. Consider a set of pairwise images $\{S^{n},T^{n}\}_{n=1}^{N}$, we are now ready to define the joint loss of GDN as
$L(\theta_{r},\theta_{v})$ $=\sum_{n=1}^{N}\lambda\,\|(S^{n}(\phi^{n}_{\tau}(\theta_{v}))-T^{n}\|^{2}_{2}+\frac{1}{2}(\mathcal{L}v^{n}_{0}(\theta_{v}),v^{n}_{0}(\theta_{v}))$
@@
where $\lambda$ and $\eta$ are positive weighting parameters to balance the image matching term and the geodesic loss defined in Eq. (4).
@@ 4.1 Datasets
We demonstrate our model on three synthetic datasets: $2$D synthetic circles with random radius, $2$D hand-written digits MNIST (LeCun, 1998), and Google QuickDraw (Jongejan et al., 2016) - a collection of categorized drawings contributed by online players in a drawing game. For the QuickDraw dataset, we randomely choose $10000$ images that include different classes of envelopes, moons, triangles, shirts, and etc. All images were upsampled to the size of $64\times 64$, and were pre-aligned with affine transformations within each class.
3D brain MRI. We include $828$ T1-weighted 3D brain MRI scans from the Open Access Series of Imaging Studies (OASIS-3) dataset (Fotenos et al., 2005). This dataset is part of the Learn2Reg challenge (Alsinan et al., 2022) and has been widely used in the literature as a standard benchmark (Yang et al., 2017; Wang and Zhang, 2020; Hinkle et al., 2018; Balakrishnan et al., 2019; Chen et al., 2022). The MRI scans were resampled to the dimension of $128\times 128\times 128$, with an isotropic resolution of $1.25\text{mm}^{3}$. All MRIs have undergone skull-stripping, intensity normalization, bias field correction, and affine alignment. Due to the difficulty of preserving the diffeomorphic property across individual subjects particularly with large age variations, we carefully evaluate images from subjects aged from $60$ to $90$. We split the dataset into $70\%$ for training, $15\%$ for validation, and $15\%$ for testing. The testing volumes include manually delineated anatomical structures, such as White-Matter (WM), Cerebral-Cortex (CerebralC), Ventricle (Ven), Cerebellum-Cortex (CerebellumC), Thalamus (Tha), Putamen (Puta), Caudate (Caud), Hippocampus (Hipp), and brain stem (Stem).
@@ 4.2 Experimental Design
We evaluate the effectiveness of our proposed GDN in learning geodesic deformations on 2D synthetic data. Specifically, we randomly select $12,000$ image pairs from a combined dataset of MNIST ($2,000$ per digit ranging from 00 to $4$) and binary circles ($2,000$ pairs). The dataset is partitioned into $70\%$ for training, $15\%$ for validation, and $15\%$ for testing. We compare the predicted velocity fields, $\{v_{t}\}$, and their associated transformations, $\{\phi_{t}\}$, with two optimization-based algorithms - (i) the numerical solutions derived from the original geodesic shooting equation (Eq. (2)) and (ii) predictions from NeurEPDiff (Wu and Zhang, 2023). We also visualize a sequence of deformed images along the predicted geodesic path.
@@
Evaluation on generalizability and regularizability. To investigate the generalizability and regularizability of GDN in the context of registration, we first train the model on 2D circles and 2D MNIST data (digits 0-4 only) and test it on data from the same dataset but was not seen during training. We then evaluate GDN’s performance across different distributions of training and testing data. This includes testing the trained model on the remaining digits (digits 5-9) and a distinct dataset from Google QuickDraw. The final deformed images vs. target images are produced to examine the quality of image alignment. Meanwhile, we visualize the predicted deformations and their determinant of Jacobian maps to examine whether they are well regularized. Additionally, we report the negative values observed in the determinant of Jacobian maps across all methods. We compare the performance of GDN with the state-of-the-art deep learning based diffeomorphic registration methods, notably LagoMorph (LM) (Hinkle et al., 2018), QuickSilver (QS) (Yang et al., 2017), VoxelMorph-diff (VM) (Dalca et al., 2019), and TransMorph (TM) (Chen et al., 2022). Please note that only the LM and QS methods are based on LDDMM. The other two baseline approaches are based on stationary velocity fields and do not account for the geodesic path. All methods are trained with their best performance reported.
Final deformation evaluation. We evaluate the quality of predicted deformations of GDN by quantitatively measuring its accuracy on real 3D brain MRIs. We perform registration-based segmentation and compare the resulting segmentation accuracy of our GDN with all baseline algorithms using two distinct quantification metrics. One metric is the dice similarity coefficient (DSC) (Dice, 1945) that evaluates volume overlap between the propagated/deformed segmentation $A$ and the manual segmentation $B$ for each structure. Such a dice score can be computed by DSC$(A,B)=2(|A|\cap|B|)/(|A|+|B|)$, where $\cap$ denotes an intersection of two regions. Another is the Hausdorff distance (HD) (Huttenlocher et al., 1993), which measures the maximum discrepancy between boundaries of two segmented labels $A$ and $B$. Given two sets of boundary points $X\in A$ and $Y\in B$ for a specific biological structure, we compute the distance score by HD$(X,Y)=max\{h(X,Y),h(Y,X)\}$, where $h(X,Y)=\mathop{max}_{x_{i}\in X}\mathop{min}_{y_{i}\in Y}\|x_{i}-y_{i}\|$ and vice versa.
Parameter setting. The training of all our experiments is implemented on an server with AMD EPYC $7502$ CPU of $126$GB memory and Nvidia GTX $3090$Ti GPUs. We train $2$D and $3$D networks using the Adam optimizer (Kingma and Ba, 2014) with weight decay as $1e^{-4}$. The $2$D network was trained for $1000$ epochs with a batch size of $100$, while the $3$D network was trained for $2000$ epochs with a batch size of $4$. We used a learning rate of $1e^{-3}$ for our network GDN. We conduct a grid search over a range of registration parameters $(\sigma,\gamma,\alpha)$, taking into account their mutual influence on one another, and found the combination of $(0.03,1.0,0.5)$ for $2$D experiments and $(0.03,2.0,0.1)$ for $3$D experiments yield the best results.
@@
The numerical solution of the geodesic shooting equation (Eq. (2)) that guides the learning of GNO is generated in real-time during the training process. In all experiments, we opt for a commonly employed Euler integrator to numerically solve this shooting equations (Yang et al., 2017; Wang and Zhang, 2020; Hinkle et al., 2018; Vialard et al., 2012b; Beg et al., 2005). Other advanced integrators such as Runge–Kutta can be easily applied. We set all integration steps as $10$, the smoothing parameters for the operator $K$ in Eq. (2) as $\alpha=3.0,c=3.0$, the positive weighting parameter of registration as $\lambda=0.03$. The number of hidden layers in GNO is set as $T=4$.
@@ 4.3 Experimental Results
In Fig. 2, we present the visualizations of the predicted transformations from GDN alongside numerical solutions obtained from two optimization-based algorithms, the original geodesic shooting equation and NeurEPDiff. The deformation processes generated by both GDN and numerical solution show a nice degree of similarity, while GDN slightly outperforms NeurEPDiff. Furthermore, we compute quantitative metrics (MSE error) of deformed images, transformation fields, and velocity fields between the real numerical solution and our method at each time integration step (as shown in Fig. 3). These findings collectively indicate that GDN effectively learns geodesic mapping functions, demonstrating competitive quality in comparison to real numerical solutions.
Figure 2: Visulization of predicted geodesics from GDN vs. real numerical solutions (Zhang and Fletcher, 2015) and predictions from NeurEPDiff (Wu and Zhang, 2023). Left to right: source and target images, predicted geodesic deformations along time $t$. Top to bottom: deformed images, transformation fields, and velocity fields).
Fig. 4 visualizes examples of predicted transformation fields and their associated determinant of Jacobian (DetJac) maps on in-distribution (ID) testing data from all methods. Different values of DetJac indicate different patterns of volume changes. For example, a DetJac value of $1$ indicates no volume change, while DetJac $<1$ reflects volume shrinkage and DetJac $>1$ implies volume expansion. A DetJac value smaller than zero suggests an artifact or singularity in the transformation field, i.e., a failure to regularize the smoothness of deformation fields when the effect of folding and crossing grids occurs. While the value of DetJac from all methods do not suggest violations of regularization, our work GDN exhibits smoother transformation grids, indicating more effective regularization. In particular, the estimated transformations from GDN are well regularized to concentrated primarily on the shape of circles or digits. Deformation changes on background areas (particularly with no intensity changes between images) have little-to-zero values.
@@
Figure 3: A comparison of predicted geodesics by GDN vs. numerical integration of EPDiff. Left to right: a error (MSE) between deformed images, transformations, and velocities along the geodesic path.
@@
Figure 4: A comparison of predicted transformation grids and their associated determinant of Jacobian maps on in-distribution (ID) testing data from our method GDN and other baselines.
Figure 5: A comparison of registration results on out-of-distribution (OOD) testing data from all methods: GDN, VoxelMorph-diff (VM-diff), LagoMorph (LM), and Quicksilver (QS). Left to right: source and target images, deformed source images, deformation fields, and error maps of deformed source vs. target images.
The top panel of Fig. 6 presents visual examples of source, target, and deformed images derived from real 3D brain MRIs for all methods, showcasing axial, sagittal, and coronal views. The bottom panel of Fig. 6 presents groundtruth segmentation maps overlaid with images vs. propagated segmentation maps generated by GDN and all baseline algorithms.
@@
Figure 6: Top to bottom: examples of 3D brain MRIs (in coronal, axial, and sagittal views) and registration-based segmentation maps overlaid on brain MRIs. Left to right: source and target images; deformed source images across all methods.
@@
Fig. 7 displays a quantitative comparison between the manually labeled vs. propagated segmentation labels deformed by deformations fields predicted from all methods. The top panel reports the statistics of dice scores (the higher the better), and the bottom panel reports the statistics of Hausdorff distance (the lower the better) over eight brain structures from hundreds of testing registration pairs. It shows that our method GDN is able to produce superior quality of final deformation fields in comparison with other baselines. Overall, our GDN achieves notably higher dice and lower Hausdorff distance on average. Meanwhile, an additional benefits of GDN is to forgo the computation of geodesic shooting equation inherent in conventional LDDMM-based registration networks. This step can substantially increase computational cost in the testing phase, particularly when high-dimensional images. In this experiment, GDN demonstrates its capability to predict geodesics in $0.013s$, outperforming the other two baselines in the context of LDDMM - LM and QS ($0.025$ seconds). It’s important to note that the other two baselines - VM-diff (Balakrishnan et al., 2019) and TM (Chen et al., 2022) models employ stationary velocity fields to parameterize the deformation fields, treating the velocities as constant over time. However, these methods are not guaranteed to provide a geodesic in the deformation spaces.
@@
Figure 7: Top to bottom: a comparison of dice score vs. Hausdorff distance by propagating the deformation field to the segmentation labels on nine brain structures White-Matter (WM), Cerebral-Cortex (CerebralC), Ventricle (Ven), Cerebellum-Cortex (CerebellumC), Thalamus (Tha), Putamen (Puta), Caudate (Caud), Hippocampus (Hipp), and brain stem (Stem).
@@
Fig. 8 presents a quantitative comparison of negative values in the determinant of Jacobian(DetJac) of the transformations across three different datasets — in-distribution (ID) testing data, out-of-distribution (OOD) testing data, and real 3D brain MRIs. The results indicate that for both the ID testing data and the 3D brain MRIs, there are barely violations of regularization. It is worth noting that our method, GDN, demonstrates advantages on the OOD testing data, with a zero number of negative values reported. These findings indicate the potential of GDN achieving improved generalizability and regularization compared to other baselines.
@@
Figure 8: A report on the negative values in determinant of Jacobian of the transformations on three different datasets — in-distribution (ID) testing data, out-f-distribution (OOD) testing data, and real 3D brain MRIs.

### v3

@@ 3.2 Joint Learning of GDN
The loss function of GDN integrates losses from both an unsupervised registration network and the geodesic neural operators informed by real numerical solutions. Given a set of pairwise images $\{S^{n},T^{n}\}_{n=1}^{N}$, we are now ready to define the joint loss of GDN as
$L(\theta_{r},\theta_{v})$ $=\sum_{n=1}^{N}\lambda\,\|(S^{n}(\phi^{n}_{\tau}(\theta_{v}))-T^{n}\|^{2}_{2}$
$+\frac{1}{2}\xi(\mathcal{L}v^{n}_{0}(\theta_{v}),v^{n}_{0}(\theta_{v}))$
@@
where $\lambda$, $\xi$, and $\eta$ are positive weighting parameters to balance the image matching term and the geodesic loss defined in Eq. (3.1).
@@ 4.1 Datasets
We demonstrate our model on three synthetic datasets: 22D synthetic circles with random radius, 22D hand-written digits MNIST (LeCun, 1998), and Google QuickDraw (Jongejan et al., 2016) - a collection of categorized drawings contributed by online players in a drawing game. For the QuickDraw dataset, we randomly choose $10000$ images that include different classes of envelopes, moons, triangles, shirts, etc. All images were upsampled to the size of $64\times 64$, and were pre-aligned with affine transformations within each class.
3D brain MRI. We include $828$ T1-weighted 3D brain MRI scans from the Open Access Series of Imaging Studies (OASIS-3) dataset (Fotenos et al., 2005). This dataset is part of the Learn2Reg challenge (Alsinan et al., 2022) and has been widely used in the literature as a standard benchmark (Yang et al., 2017; Wang and Zhang, 2020; Hinkle et al., 2018; Balakrishnan et al., 2019; Chen et al., 2022). The MRI scans were resampled to the dimension of $128\times 128\times 128$, with an isotropic resolution of $1.25\text{mm}^{3}$. All MRIs have undergone skull-stripping, intensity normalization, bias field correction, and affine alignment. Due to the difficulty of preserving the diffeomorphic property across individual subjects, particularly with large age variations, we carefully evaluate images from subjects aged $60$ to $90$. We split the dataset into $70\%$ for training, $15\%$ for validation, and $15\%$ for testing. The testing volumes include manually delineated anatomical structures, such as White-Matter (WM), Cerebral-Cortex (CerebralC), Ventricle (Ven), Cerebellum-Cortex (CerebellumC), Thalamus (Tha), Putamen (Puta), Caudate (Caud), Hippocampus (Hipp), and brain stem (Stem).
@@ 4.2 Experimental Design
We evaluate the effectiveness of our proposed GDN in learning geodesic deformations on 2D synthetic data. Specifically, we randomly select $12,000$ image pairs from a combined dataset of MNIST ($2,000$ per digit ranging from 0 to $4$) and binary circles ($2,000$ pairs). The dataset is partitioned into $70\%$ for training, $15\%$ for validation, and $15\%$ for testing. We compare the predicted velocity fields, $\{v_{t}\}$, and their associated transformations, $\{\phi_{t}\}$, with two optimization-based algorithms - (i) the numerical solutions derived from the original geodesic shooting equation (Eq. (2)) and (ii) predictions from NeurEPDiff (Wu and Zhang, 2023). We also visualize a sequence of deformed images along the predicted geodesic path.
@@
Evaluation on generalizability and regularizability. To investigate the generalizability and regularizability of GDN in the context of registration, we first train the model on 2D circles and 2D MNIST data (digits 0-4 only) and test it on data from the same dataset but were not seen during training. We then evaluate GDN’s performance across different distributions of training and testing data. This includes testing the trained model on the remaining digits (digits 5-9) and a distinct dataset from Google QuickDraw. The final deformed images vs. target images are produced to examine the quality of image alignment. Meanwhile, we visualize the predicted deformations and their determinant of Jacobian maps to examine whether they are well regularized. Additionally, we report the percentage of voxels with a negative Jacobian determinant on the 2D dataset and the negative values observed in the determinant of Jacobian maps across all methods for real 3D brain MRI datasets. We compare the performance of GDN with the state-of-the-art deep learning-based diffeomorphic registration methods, notably LagoMorph (LM) (Hinkle et al., 2018), QuickSilver (QS) (Yang et al., 2017), VoxelMorph-diff (VM) (Dalca et al., 2019), and TransMorph (TM) (Chen et al., 2022). Please note that only the LM and QS methods are based on LDDMM. The other two baseline approaches are based on stationary velocity fields and do not necessarily generate geodesic paths unless the space is equipped with a flat affine connections (Lorenzi and Pennec, 2013). All methods are trained with their best performance reported.
[…]
Statistical evaluation. We performed paired t-tests (Student, 1908) to compare our proposed model GDN with other benchmarks and reported the corresponding p-values for all models on the DSC scores across different anatomical structures using real 3D brain MRI datasets. This analysis provides insight into whether GDN performs significantly better or differently compared to the baselines in terms of registration accuracy.
Parameter setting. The training of all our experiments is implemented on an server with AMD EPYC $7502$ CPU of 126126GB memory and Nvidia GTX 30903090Ti GPUs. We train 22D and 33D networks using the Adam optimizer (Kingma and Ba, 2014) with weight decay as $1e^{-4}$. The 22D network was trained for $1000$ epochs with a batch size of $100$, while the 33D network was trained for $2000$ epochs with a batch size of $6$. We used a learning rate of $5e^{-4}$ for our network GDN. We set the number of integration steps to $\tau=10$ for the 2D dataset and $\tau=5$ for the 3D dataset to balance accuracy and efficiency. The smoothing parameters for the operator $K$ in Eq. (2) are set as $\alpha=1.0$, $\beta=0.5$ for the 2D dataset, and $\alpha=1.0$, $\beta=1.0$ for the 3D dataset. The positive weighting parameter for registration is $\lambda=\frac{1}{{0.03}^{2}}$, with the regularization weight set to $\xi=2.0$ for 2D and $\xi=0.15$ for 3D. The number of hidden layers in the GNO is configured as $J=4$ for 2D and $J=1$ for 3D. The backbone of the registration network is a U-Net architecture with a kernel size of 3. For our GDN, the channel configurations are set to [[8, 16, 16, 8], [8, 16, 16, 16, 8]]. All baseline models, including LM, QS, and VM-diff, use the same U-Net architecture with identical kernel size and channel configurations to ensure a fair comparison.
Ablation Study. We conduct an ablation study on 3D real brain MRIs to evaluate the influence of the number of integration steps ($\tau$) in the geodesic learning module on both registration accuracy and regularity.
Please note that, our model GDN, when excluding the proposed GNO module and geodesic loss, is equivalent to the baseline LagoMorph. This serves as an ablation study to assess the contribution of the GNO and geodesic loss components.
@@
The numerical solution of the geodesic shooting equation (Eq. (2)) that guides the learning of GNO is generated in real-time during the training process. In all experiments, we opt for a commonly employed Euler integrator to numerically solve this shooting equation (Yang et al., 2017; Wang and Zhang, 2020; Hinkle et al., 2018; Vialard et al., 2012b; Beg et al., 2005). Other advanced integrators, such as Runge-Kutta, can be easily applied.
@@ 4.3 Experimental Results
Figure 2: Visulization of predicted geodesics from GDN vs. real numerical solutions (Zhang and Fletcher, 2015) and predictions from NeurEPDiff (Wu and Zhang, 2023). Left to right: source and target images, predicted geodesic deformations along time $t$. Top to bottom: deformed images, transformation fields, and velocity fields.
In Fig. 2, we present the visualizations of the predicted transformations from GDN alongside numerical solution obtained from two optimization-based algorithms, the original geodesic shooting equation and NeurEPDiff. The deformation processes generated by both GDN and numerical solution show a nice degree of similarity, while GDN slightly outperforms NeurEPDiff.
Figure 3: Top to bottom: Comparison of predicted geodesics by GDN/NeurEPDiff vs. numerical integration of EPDiff. Left to right: MSE between deformed images, transformations, and velocities along the geodesic path.
Fig. 3 compares the discrepancy between our method and the numerical solution to the EPDiff equation by using NeurEPDiff as a baseline for reference. Specifically, we quantify the mean squared error (MSE) between the predicted time-varying velocity fields and the geodesic velocities obtained through numerical integration of EPDiff. Additionally, we assess the similarity of the resulting transformations and deformed images along the predicted and ground-truth geodesic paths. This result shows that our method, GDN, achieves slightly better performance than NeurEPDiff across these metrics.
Fig. 4 visualizes examples of predicted transformation fields and their associated determinant of Jacobian (DetJac) maps on in-distribution (ID) testing data from all methods. Different values of DetJac indicate different patterns of volume changes. For example, a DetJac value of $1$ indicates no volume change, while DetJac <1<1 reflects volume shrinkage and DetJac >1>1 implies volume expansion. A DetJac value smaller than zero suggests an artifact or singularity in the transformation field, i.e., a failure to regularize the smoothness of deformation fields when the effect of folding and crossing grids occurs. While the value of DetJac from all methods do not suggest violations of regularization, our work GDN exhibits smoother transformation grids, indicating more effective regularization. In particular, the estimated transformations from GDN are well regularized to concentrated primarily on the shape of circles or digits. Deformation changes on background areas (particularly with no intensity changes between images) have little-to-zero values.
Figure 4: A comparison of predicted transformation grids and their associated determinant of Jacobian maps on in-distribution (ID) testing data from our method GDN vs. baselines. From top to bottom: source and target image pairs, resulting deformed images, predicted transformation grids, and determinant of Jacobian maps.
@@
Figure 5: A comparison of predicted transformation grids and their associated determinant of Jacobian maps on out-of-distribution (OOD) testing data from our method GDN vs. baselines. Top to bottom: source and target images, deformed source images, predicted deformation fields, and determinant of Jacobian maps.
Fig. 6 presents visual comparisons of deformed images, registration-based segmentation maps overlaid on 3D brain MRIs, predicted deformation fields, and determinants of Jacobian maps for all methods, shown in both sagittal and coronal views. Our method demonstrates a similar smoothing effect on transformation grids with LM and QS, while more clearly highlighting biologically plausible growth patterns, such as ventricular enlargement (highlighted in red in the determinants of Jacobian maps). These patterns are more interpretable than those produced by the baseline methods, VM and TM.
@@
Figure 6: Top to bottom: examples of 3D brain MRIs (demonstrated in coronal and sagittal views), registration-based segmentation maps overlaid on brain MRIs, predicted deformation fields, and determinants of Jacobian maps. Left to right: source and target images; deformed source images (segmentation labels) across all methods.
@@
Fig. 7 displays a quantitative comparison between the manually labeled vs. propagated segmentation labels deformed by deformations fields predicted from all methods. The top panel reports the statistics of dice scores (the higher the better) over eight brain structures from hundreds of testing registration pairs. The bottom panel summarizes the corresponding means and standard deviations. It shows that our method GDN is able to produce superior-quality final deformation fields in comparison with other baselines. Overall, our GDN achieves notably higher Dice on average. Meanwhile, an additional benefit of GDN is to forgo the computation of the geodesic shooting equation inherent in conventional LDDMM-based registration networks. This step can substantially increase computational cost in the testing phase, particularly when high-dimensional images. In this experiment, GDN demonstrates its capability to predict geodesics in $0.026s$, outperforming the other two baselines in the context of LDDMM - LM and QS ($0.036$ seconds). It’s important to note that the other two baselines - VM-diff (Balakrishnan et al., 2019) and TM (Chen et al., 2022) models employ stationary velocity fields to parameterize the deformation fields, treating the velocities as constant over time.
@@
Figure 7: Top: A comparison of dice score by propagating the deformation field to the segmentation labels on nine brain structures White-Matter (WM), Cerebral-Cortex (CerebralC), Ventricle (Ven), Cerebellum-Cortex (CerebellumC), Thalamus (Tha), Putamen (Puta), Caudate (Caud), Hippocampus (Hipp), and brain stem (Stem).
We also report the corresponding p-values from pairwise statistical tests comparing our proposed model, GDN, with four baseline methods across the anatomical brain structures based on Dice scores. The p-values for GDN vs. LM (<0.001<0.001), QS (<0.001<0.001), VM (<0.001<0.001), and TM (<0.001<0.001) are all well below the significance threshold of $0.05$, indicating that GDN performs significantly differently compared with baselines.
@@
Fig. 8 presents a quantitative comparison of negative values in the determinant of Jacobian(DetJac) of the transformations across three different datasets — in-distribution (ID) testing data, out-of-distribution (OOD) testing data, and real 3D brain MRIs. The results indicate that for both the ID testing data and the 3D brain MRIs, there are barely violations of regularization for all methods. However, it is worth noting that our method, GDN, demonstrates advantages on the OOD testing data, with zero negative values reported. However, when tested on data outside the training distribution, the percentage of negative Jacobian determinants begins to diverge across models: GDN (0.005%), LM (0.132%), QS (0.255%), VM-Diff (0.038%), and TM (0.011%). These findings indicate the potential of GDN to achieve improved generalizability and regularization compared to other baselines.
@@
Figure 8: A report on the negative values in determinant of Jacobian of the transformations on three different datasets: in-distribution (ID) testing data, out-of-distribution (OOD) testing data, and real 3D brain MRIs.
LDDMM-Based SVF-Based
GDN LM QS VM-Diff TM
Number of Parameters 79.3K 56.5K 56.5K 56.5K 46.5M
Training (per epoch) 61.5s 14.61s 20.5s 20.78s 33.39s
Testing 0.030s 0.043s 0.043s 0.017s 0.020s
Table 1: A summary of model parameters, training time per epoch, and testing time on 3D brain MRI data.
Tab. 1 summarizes the model parameters, training time per epoch, and inference time for GDN versus the baselines. Our model, GDN, has a larger number of parameters than LM and QS (under the same LDDMM setting) due to the added geodesic learning module. In addition, training GDN requires solving the EPDiff equation numerically to supervise the geodesic loss. These two factors contribute to GDN’s longer training time per epoch compared to LM and QS. However, a key advantage of GDN lies in inference: once trained, GDN does not require numerical integration during testing. As a result, GDN achieves faster inference time (0.030s) than LM and QS (0.043s), and we expect this advantage to become more pronounced with higher-dimensional image data. When compared to VM-diff and TM under the stationary velocity field setting, GDN does not show a significant advantage in training or inference time, primarily due to its larger model size relative to VM-diff. The model TM, while larger in size, benefits from parallel computation and hence is more time-efficient.
Fig. 9 presents a quantitative comparison of the average dice scores on 3D brain MRIs with different integration steps of our proposed module GNO. The results show that our method achieves stable dice accuracy based on the deformed template segmentation labels as the number of integration steps increases from five to ten.
Figure 9: A comparison of average dice score on 3D brain MRIs with a varying number of integration steps.